OpenStudy (anonymous):

Solve By System Of Elimination : -5y+7y=13 -4x+14y=2

OpenStudy (anonymous):

I can help you :)

OpenStudy (anonymous):

Help Please With Steps.

OpenStudy (anonymous):

Ok :)

OpenStudy (anonymous):

1. -2(-5x+7y=13) <---distribute the -2 and tell me what u get

OpenStudy (anonymous):

I KNOW The Answer

OpenStudy (anonymous):

But The Steps I Dont Know.

OpenStudy (anonymous):

-4x+14y=2 -10x-14y=-26